# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 2.3. Global Relative Entropy Control
(×$\displaystyle\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}-\frac{1}{N}V\ast\omega(x_{1})+\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
Using the simple Cauchy-Schwartz inequality, we bound by
$\displaystyle\frac{1}{N^{2}}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}|V\ast\omega(x_{1})|^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle$++ $\displaystyle\frac{2}{N^{2}}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}|V\ast\omega(x_{1})|\Big{|}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle$++ (×$\displaystyle\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
@@
$\frac{C}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}G^{N}\,\mathrm{d}Z^{N}=\frac{C}{N}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x|^{2}}{(1+t)^{2}}\Big{)}G^{N,1}\,\mathrm{d}z,$
where we make use of the linear growth of $\nabla\log g$ in view of Proposition 4.1 below. By the Gibbs inequality Lemma 2.3 with $N=1$ and $\eta=(1+t)/C$, and using the Gaussian upper bound for $g$ as in Lemma 3.2 below, we have
$\displaystyle\frac{C}{N}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x|^{2}}{(1+t)^{2}}\Big{)}G^{N,1}\,\mathrm{d}z$
@@
(×$\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle$++ (×$\displaystyle\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1})-{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
In the following two steps we shall deal with $I_{1}$ and $I_{2}$ respectively.
Step 3: We first give an estimate for $I_{1}$ using our logarithmic gradient estimate Proposition 4.1 and the extended version of the Law of Large Numbers Theorem 2.3. Notice that for $|x_{1}|\leq 2|x_{j}|$ we have
$\displaystyle|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})||\nabla\log g(m_{1},x_{1})|$
@@
due to the boundedness of $V$ and the linear growth of $\nabla\log g$ in view of Proposition 4.1 below, while for $|x_{1}|\geq 2|x_{j}|$ we have
$\displaystyle|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})||\nabla\log g(m_{1},x_{1})|$
@@
+|×$\displaystyle+\Big{|}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m\Big{(}V(x_{1})-V(x_{1}-y)\Big{)}g(m,y)\,\mathrm{d}m\,\mathrm{d}y\Big{|}\Big{(}\frac{\sqrt{1+\log(1+t)}}{\sqrt{1+t}}+\frac{|x_{1}|}{1+t}\Big{)}$
@@
$\displaystyle+C\int_{|x_{1}|\geq 2|y|}m|y|g(m,y)\sup_{s\in[0,1]}|\nabla V(x_{1}-sy)|\Big{(}\frac{\sqrt{1+\log(1+t)}}{\sqrt{1+t}}+\frac{|x_{1}|}{1+t}\Big{)}\,\mathrm{d}m\,\mathrm{d}y$
[…]
$\displaystyle+C\int_{|x_{1}|\geq 2|y|}mg(m,y)\Big{(}\frac{\sqrt{1+\log(1+t)}}{\sqrt{1+t}}\frac{|y|}{|x_{1}|}+\frac{|y|}{1+t}\Big{)}\,\mathrm{d}m\,\mathrm{d}y$
@@
and the Gaussian upper bound of $g(z)$ as in Lemma 3.2 below. Hence we conclude that
$I_{1}=\frac{1}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{|}\frac{1}{\sqrt{N}}\sum_{j=2}^{N}\psi(z_{1},z_{j})\Big{|}^{2}G^{N}\,\mathrm{d}Z^{N},$
@@
$\sup_{x_{1}}|\psi(x_{1},x_{j})|\leq C\Big{(}\frac{\sqrt{1+\log(1+t)}}{\sqrt{1+t}}+\frac{|x_{j}|}{1+t}\Big{)}.$
It is also direct to check that $\psi$ satisfies the cancelling condition (2.4). Now we conclude with the Gibbs inequality Lemma 2.3 and the extended version of the Law of Large Numbers Theorem 2.3 with some $\eta(t)=C^{-1}\frac{\sqrt{1+t}}{\sqrt{1+\log(1+t)}}$:
$\displaystyle I_{1}$ $\displaystyle\leq\frac{1}{\eta^{2}}H_{N}(G^{N}|g^{N})+\frac{1}{\eta^{2}N}\log\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\exp\Big{(}\frac{\eta^{2}}{N}\sum_{j,k=1}^{N}\psi(z_{1},z_{j})\psi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}$
$\displaystyle\leq C\frac{1+\log(1+t)}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
Step 4: We now give an estimate for $I_{2}$ using its special separation structure, namely
$I_{2}\leq\frac{C}{N^{2}}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}G^{N}\,\mathrm{d}Z^{N}.$
We shall first integrate with respect to the variable $z_{1}$, since now the term in the summation is no longer dependent on $x_{j}$. Denote by $P_{z_{2},...,z_{N}}^{N,1}(z_{1})$ the conditional probability measure of $z_{1}$ under $(z_{2},...,z_{N})$, i.e.
$P_{z_{2},..,z_{N}}^{N,1}(z_{1})=\mathbb{P}(Z_{1}|Z_{2},...,Z_{N}).$
@@
$P_{z_{2},..,z_{N}}^{N,1}(z_{1})=\frac{G^{N}(z_{1},...,z_{N})}{G^{N,N-1}(z_{2},...,z_{N})}.$
Now we are able to rewrite $I_{2}$ as
$I_{2}\leq\frac{C}{N^{2}}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N-1}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}G^{N,N-1}\,\mathrm{d}Z^{N-1}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}P_{z_{2},..,z_{N}}^{N,1}(z_{1})\,\mathrm{d}Z.$
Using our Gibbs inequality Lemma 2.3 for one variable with some $\eta(t)=C^{-1}(1+t)/(1+\log(1+t))$, we have
×$\displaystyle\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}P_{z_{2},..,z_{N}}^{N,1}(z_{1})\,\mathrm{d}Z$
@@
Hence our estimate for $I_{2}$ writes
$\displaystyle I_{2}\leq$ $\displaystyle\,\frac{C}{N^{2}}\frac{1+\log(1+t)}{1+t}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N-1}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}G^{N,N-1}\,\mathrm{d}Z^{N-1}$
$\displaystyle+C\frac{1+\log(1+t)}{1+t}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N-1}}H_{1}(P_{z_{2},..,z_{N}}^{N,1}|g)G^{N,N-1}\,\mathrm{d}Z^{N-1}.$
The first integral is $O(1/N)$ due to the time-independence of ${\mathcal{M}}^{N}$ and the fact that $G_{0}^{N,N-1}=g_{0}^{\otimes(N-1)}$ (one can also verify by another application of the Gibbs inequality Lemma 2.3 and the Law of Large Numbers Theorem 2.1, since all terms in the summation are bounded; this will also work for the non-tensorized initial data case), while the second integral is controlled by $CH_{N}(G^{N}|g^{N})$ using the chain rule of relative entropy Lemma 2.4, with
$K_{x}^{1}=P_{z_{2},...,z_{N}}^{N,1},\quad K_{x}^{2}=g,\quad m_{1}=G^{N,N-1},\quad m_{2}=g^{N-1}.$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{N}(G^{N}|g^{N})\leq C\frac{1+\log(1+t)}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
The Grönwall’s inequality will give our desired bound for the global relative entropy. Applying the sub-additivity property and the Csiszár-Kullback-Pinsker inequality as in [SZ24], we resolve the $L^{1}$ strong propagation of chaos.
@@
To close this subsection, we make a short sketch of the proof of the global relative entropy control on the torus case, without the independence assumption on ${\mathcal{M}}$ and $X(0)$. One is easy to observe that the time evolution of the relative entropy is invariant in this case, once we change the integration domain into the torus. Hence we only need to bound the two terms (2.8) and (2.10).
@@
which is the same as [SZ24], then it is direct to show that $\nabla\log g$ and $\nabla^{2}\log g$, and thus the test functions in (2.8) and (2.10), are uniformly bounded from above by some universal constant. Hence this is essentially the same condition with [SZ24]. Moreover, by applying the time-uniform LSI on the limit density, as pointed out in [GLBM24], one can even obtain uniform-in-time propagation of chaos results for this torus model.

### v2

@@ 2.3. Global Relative Entropy Control
$\displaystyle\int_{\mathbb{D}^{N}}G^{N}\Big{|}-\frac{1}{N}V\ast\omega(x_{1})+\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
Using the simple Cauchy-Schwarz inequality, we bound by
$\displaystyle\frac{1}{N^{2}}\int_{\mathbb{D}^{N}}G^{N}|V\ast\omega(x_{1})|^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle$++ $\displaystyle\frac{2}{N^{2}}\int_{\mathbb{D}^{N}}G^{N}|V\ast\omega(x_{1})|\Big{|}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle$++ $\displaystyle\int_{\mathbb{D}^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
@@
$\frac{C}{N}\int_{\mathbb{D}^{N}}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}=\frac{C}{N}\int_{\mathbb{D}}|\nabla\log g(m_{1},x_{1})|^{2}G^{N,1}\,\mathrm{d}z_{1}.$
Applying the Donsker-Varadhan inequality in Lemma 2.3 with $N=1$ and $\eta=(1+t)/(2B_{1})$, and using the Gaussian upper bound for $g$ as in Lemma 3.2 below, we have
$\displaystyle\frac{C}{N}\int_{\mathbb{D}}|\nabla\log g(m_{1},x_{1})|^{2}G^{N,1}\,\mathrm{d}z_{1}$
[…]
@@
$\int_{\mathbb{D}^{N}}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle$++ $\displaystyle\,\int_{\mathbb{D}^{N}}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1})-{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
In the following two steps, we shall deal with $J_{1}$ and $J_{2}$ respectively.
Step 3: We first give an estimate for $J_{1}$ using our logarithmic gradient estimate Proposition 4.1 and the extended version of the Law of Large Numbers Theorem 2.3. We need to prove that the test function in $J_{1}$ is uniformly bounded in $z_{1}$.
For $|x_{1}|\leq 2|x_{j}|$ we have
$|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})|^{2}\leq C$
by trivial bounds. Consequently,
$\displaystyle|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})|^{2}|\nabla\log g(m_{1},x_{1})|^{2}$
@@
thanks to the logarithmic gradient estimate in Proposition 4.1 below.
For $|x_{1}|\geq 2|x_{j}|$ we divide the expression in the large bracket in $J_{1}$ into two parts. For the first part, we have
$\displaystyle|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})|\leq C|x_{j}|\sup_{s\in[0,1]}|\nabla V(x_{1}-sv_{j})|\leq C\frac{|x_{j}|}{|x_{1}|},$
where we use the mean-value inequality and that
$\sup_{s\in[0,1]}|\nabla V(x_{1}-sy)|\leq\frac{C}{|x_{1}|}\text{ for any }|x_{1}|\geq 2|y|.$
For the second part, we have
${\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})=\int_{\mathbb{D}}m\Big{(}V(x_{1})-V(x_{1}-y)\Big{)}g(m,y)\,\mathrm{d}m\,\mathrm{d}y,$
which is controlled by
$\displaystyle\,C\int_{|x_{1}|\leq 2|y|}mg(m,y)\,\mathrm{d}m\,\mathrm{d}y+C\int_{|x_{1}|\geq 2|y|}m|y|g(m,y)\sup_{s\in[0,1]}|\nabla V(x_{1}-sy)|\,\mathrm{d}m\,\mathrm{d}y$
@@
and the Gaussian upper bound of $g(z)$ as in Lemma 3.2 below. This will deduce that for $|x_{1}|\geq 2|x_{j}|$:
$|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})|^{2}\leq C\Big{(}1+\frac{|x_{j}|^{2}}{|x_{1}|^{2}}+\frac{1+t}{|x_{1}|^{2}}\Big{)}.$
Consequently,
$\displaystyle|m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})|^{2}|\nabla\log g(m_{1},x_{1})|^{2}$
[…]
To this end, we conclude that
$J_{1}=\frac{1}{N}\int_{\mathbb{D}^{N}}\Big{|}\frac{1}{\sqrt{N}}\sum_{j=2}^{N}\psi(z_{1},z_{j})\Big{|}^{2}G^{N}\,\mathrm{d}Z^{N},$
@@
$\sup_{z_{1}}|\psi(z_{1},z_{j})|\leq C\Big{(}\frac{1}{\sqrt{1+t}}+\frac{|x_{j}|}{1+t}\Big{)}.$
It is also direct to check that $\psi$ satisfies the cancelling condition (2.4). Now we conclude with the Donsker-Varadhan inequality Lemma 2.3 and the extended version of the Law of Large Numbers Theorem 2.3 with some $\eta(t)=C^{-1}\sqrt{1+t}$:
$\displaystyle J_{1}$ $\displaystyle\leq\frac{1}{\eta^{2}}H_{N}(G^{N}|g^{N})+\frac{1}{\eta^{2}N}\log\int_{\mathbb{D}^{N}}\exp\Big{(}\frac{\eta^{2}}{N}\sum_{j,k=1}^{N}\psi(z_{1},z_{j})\psi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}$
$\displaystyle\leq\frac{C}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
Step 4: We now give an estimate for $J_{2}$ using its special separation structure, namely
$J_{2}\leq\frac{C}{N^{2}}\int_{\mathbb{D}^{N}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}\Big{(}\frac{1}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}G^{N}\,\mathrm{d}Z^{N}.$
We shall first integrate with respect to the variable $z_{1}$, since now the term in the summation is no longer dependent on $x_{j}$. Denote by $P_{z_{2},\cdots,z_{N}}^{N,1}(z_{1})$ the conditional probability measure of $z_{1}$ under $(z_{2},\cdots,z_{N})$, i.e.
$P_{z_{2},\cdots,z_{N}}^{N,1}(z_{1})=\mathbb{P}(Z_{1}|Z_{2},\cdots,Z_{N}).$
@@
$P_{z_{2},\cdots,z_{N}}^{N,1}(z_{1})=\frac{G^{N}(z_{1},\cdots,z_{N})}{G^{N,N-1}(z_{2},\cdots,z_{N})}.$
Now we are able to rewrite $J_{2}$ as
$J_{2}\leq\frac{C}{N^{2}}\int_{\mathbb{D}^{N-1}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}G^{N,N-1}\,\mathrm{d}Z^{N-1}\int_{\mathbb{D}}\Big{(}\frac{1}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}P_{z_{2},\cdots,z_{N}}^{N,1}(z_{1})\,\mathrm{d}z_{1}.$
Using our Donsker-Varadhan inequality Lemma 2.3 for one variable with some $\eta(t)=C^{-1}(1+t)$, we have
$\displaystyle\int_{\mathbb{D}}\Big{(}\frac{1}{1+t}+\frac{|x_{1}|^{2}}{(1+t)^{2}}\Big{)}P_{z_{2},\cdots,z_{N}}^{N,1}(z_{1})\,\mathrm{d}z_{1}$
@@
Hence our estimate for $J_{2}$ writes
$\displaystyle J_{2}\leq$ $\displaystyle\,\frac{C}{1+t}\frac{1}{N^{2}}\int_{\mathbb{D}^{N-1}}\Big{|}\sum_{j=2}^{N}(m_{j}-{\mathbb{E}}{\mathcal{M}})\Big{|}^{2}G^{N,N-1}\,\mathrm{d}z_{2}\cdots\,\mathrm{d}z_{N}$
$\displaystyle$++ $\displaystyle\,\frac{C}{1+t}\int_{\mathbb{D}^{N-1}}H_{1}(P_{z_{2},\cdots,z_{N}}^{N,1}|g)G^{N,N-1}\,\mathrm{d}z_{2}\cdots\,\mathrm{d}z_{N}.$
The first integral is $O(1/N)$ due to the time-independence of ${\mathcal{M}}^{N}$ and the fact that $G_{0}^{N,N-1}=g_{0}^{\otimes(N-1)}$ (one can also verify by another application of the Donsker-Varadhan inequality Lemma 2.3 and the Law of Large Numbers Theorem 2.1, since all terms in the summation are bounded; this will also work for the non-factorized initial data case), while the second integral is controlled by $CH_{N}(G^{N}|g^{N})$ using the chain rule of relative entropy Lemma 2.4, with
$K_{x}^{1}=P_{z_{2},\cdots,z_{N}}^{N,1},\quad K_{x}^{2}=g,\quad m_{1}=G^{N,N-1},\quad m_{2}=g^{N-1}.$
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}H_{N}(G^{N}|g^{N})\leq\frac{C}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
The Grönwall’s inequality will give our desired bound for the global relative entropy. Applying the sub-additivity property and the CKP inequality as in [52], we resolve the $L^{1}$ strong propagation of chaos.
@@
To close this subsection, we make a short sketch of the proof of the global relative entropy control on the torus case, without the independence assumption on ${\mathcal{M}}$ and $X_{0}$. It is easy to observe that the time evolution of the relative entropy is invariant in this case, once we change the integration domain into the torus. Hence we only need to bound the two terms (2.8) and (2.10).
@@
which is the same as [52], then it is direct to show that $\nabla\log g$ and $\nabla^{2}\log g$, and thus the test functions in (2.8) and (2.10), are uniformly bounded from above by some universal constant. Hence this is essentially the same condition with [52]. Moreover, by applying the time-uniform LSI on the limit density, as pointed out in [22], one can even obtain uniform-in-time propagation of chaos results for this torus model.